# Learning Unmasking Policies for Diffusion Language Models

### v2

@@ Abstract
Diffusion (Large) Language Models (dLLMs) now match the downstream performance of their autoregressive counterparts on many tasks, while holding the promise of being more efficient during inference. One particularly successful variant is masked discrete diffusion, in which a buffer filled with special mask tokens is progressively replaced with tokens sampled from the model’s vocabulary. Efficiency can be gained by unmasking several tokens in parallel, but doing too many at once risks degrading the generation quality. Thus, one critical design aspect of dLLMs is the sampling procedure that selects, at each step of the diffusion process, which tokens to replace. Indeed, recent work has found that heuristic strategies such as confidence thresholding lead to both higher quality and token throughput compared to random unmasking. However, such heuristics have downsides: they require manual tuning, and we observe that their performance degrades with larger buffer sizes. In this work, we instead propose to train sampling procedures using reinforcement learning. Specifically, we formalize masked diffusion sampling as a Markov decision process in which the dLLM serves as the environment, and propose a lightweight policy architecture based on a single-layer transformer that maps dLLM token confidences to unmasking decisions. Our experiments show that these trained policies match the performance of state-of-the-art heuristics when combined with semi-autoregressive generation, while outperforming them in the full diffusion setting. We also examine the transferability of these policies, finding that they can generalize to new underlying dLLMs and longer sequence lengths. However, we also observe that their performance degrades when applied to out-of-domain data, and that fine-grained tuning of the accuracy-efficiency trade-off can be challenging with our approach.
@@ 1 Introduction
Discrete diffusion (austin2021structured; hoogeboom2021argmax; DBLP:journals/corr/abs-2211-15089; DBLP:conf/icml/LouME24; DBLP:conf/nips/ShiHWDT24) has recently emerged as a compelling alternative to the predominant autoregressive (AR) modeling paradigm in large language models (LLMs). Unlike AR models, which generate the next token in a left-to-right fashion (radford2018improving), diffusion LLMs (dLLMs) generate text by learning to reverse a noising process that progressively corrupts a block of token sequences. In particular, masked diffusion models (MDMs; sahoo2024simple; ou2024your), a subclass of discrete diffusion models that use time-dependent BERT-style (devlin2019bert) masking as the forward noising process, have recently demonstrated impressive performance, with models like LLaDA (llada2025) and Dream (ye2025dream) matching the performance of similarly sized autoregressive LLMs.
@@
At generation time, MDMs begin with a fully masked sequence and iteratively unmask a fixed number of tokens at randomly sampled positions in each sampling step. As a result, they offer the potential for faster inference, as they can, in principle, generate multiple tokens in parallel using a single model call. Despite this promise, open-source dLLMs have, until recently, lagged behind their AR counterparts in terms of inference efficiency. This changed with Fast-dLLM (fastdllm2025), which demonstrated that a dLLM like LLaDA (llada2025) can achieve higher token throughput than similarly sized LLaMA models (dubey2024llama) while maintaining competitive performance. A key component of Fast-dLLM lies in its sampling heuristic: instead of unmasking a fixed number of randomly sampled token positions, it proposes to unmask all tokens whose confidences exceed a pre-specified threshold, resulting in adaptive sampling with a variable number of unmasked tokens per step. This success has since inspired further development of increasingly sophisticated sampling heuristics (see Section˜5 for a comprehensive overview), which has further advanced the state of the art. These heuristics can, however, be difficult to tune, and seem to work best when imposing semi-autoregressive (semi-AR) generation, in which small blocks of tokens are unmasked sequentially (arriola2025block; llada2025). We posit that such limitations are difficult to avoid with heuristics alone, since they are essentially handcrafted solutions to a sequential decision making problem: What is the optimal order in which to unmask the sequence of tokens, in order to strike a good balance between efficiency and correctness?
@@
We propose to investigate this question by moving beyond heuristics to stress instead a learning based approach, which relies on a transformer-based unmasking policy trained using reinforcement learning (RL). This approach is motivated by the above observation that unmasking in dLLMs can be viewed as a (Markovian) sequential decision making problem, and follows recent successes in reinforcement learning for post-training of language models. However, in contrast to recent works that use RL to improve the reasoning abilities of dLLMs (e.g., zhao2025d1), our goal is not to improve the capacity of the underlying model, but rather to use RL to automate the discovery of adaptive sampling strategies. We therefore treat the underlying dLLM as the environment in which to act, not the policy, leaving it unchanged and instead focusing a policy that is parameterized as a lightweight stand-alone network. Empirically, we demonstrate that our learned sampling policies match the performance of state-of-the-art heuristic samplers (fastdllm2025) in standard generation settings, while also surpassing heuristics in certain scenarios, notably when moving away from semi-AR generation. In summary, we make the following contributions:
@@
  • We conduct extensive experiments showing that, under our proposed RL framework, learned policies can match the performance of heuristic samplers such as Fast-dLLM (Section˜4.1), while also addressing some of the challenges heuristic samplers face outside the semi-AR regime (Section˜4.2).
  • We study the transferability of our learned samplers across different models and data domains (Section˜4.3), and provide detailed insights into stabilizing RL training and policy design (Section˜4.4).
@@ 2 Background
Throughout the paper, we use the notation $[L]$ to represent the set $\{1,\ldots,L\}$. We denote a sequence of tokens as ${\bm{x}}=(x^{1},\ldots,x^{d})\in{\mathcal{V}}^{d}$, where $d$ is the sequence length and ${\mathcal{V}}:=[V]$ is the vocabulary. Given our focus on masked diffusion models, we assume that the vocabulary includes a special mask token $\bm{M}$}.
@@ 2.1 Masked Diffusion Models
We focus on masked diffusion models (MDMs) (llada2025), deferring a more general introduction to discrete diffusion to Appendix E.
Training
MDMs learn to generate data by training a BERT-style (devlin2019bert) masked predictor to reverse the forward noising process. Concretely, given a training sample ${\bm{x}}_{0}\sim p_{\text{data}}$, the forward process corrupts each token independently by setting it to the mask token with probability proportional to the diffusion timestep $t\in[0,1]$:
$p_{t}({\bm{x}}_{t}\mid{\bm{x}}_{0}):=\prod_{k=1}^{d}p_{t}(x_{t}^{k}\mid{\bm{x}}_{0}),\quad\text{where}\quad p_{t}(x_{t}^{k}\mid{\bm{x}}_{0}):=\begin{cases}1-t,&\text{if }x_{t}^{k}=x_{0}^{k},\\ t,&\text{if }x_{t}^{k}={\bm{M}},\\ 0,&\text{otherwise.}\end{cases}$
Recent work has shown that an MDM parametrized by $\theta$ can learn the reverse process by maximizing the following evidence lower bound (ELBO) (ou2024your), ${\mathcal{L}}(\theta)\leq\mathbb{E}_{{\bm{x}}_{0}\sim p_{\text{data}}}[\log p_{\theta}({\bm{x}}_{0})]$:
${\mathcal{L}}(\theta):=\mathbb{E}_{t\sim U[0,1],{\bm{x}}_{0}\sim p_{\text{data}},{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t}\mid{\bm{x}}_{0})}\bigg[\frac{1}{t}\sum_{k=1}^{d}\mathbf{1}[x_{t}^{k}={\bm{M}}]\log p_{\theta}(x_{0}^{k}|{\bm{x}}_{t})\bigg]\>,$ (2.1)
with $\mathbf{1}[\cdot]$ denoting the indicator function.
Generation
When generating an answer for a given prompt $\bm{x}$}, MDMs start with a sequence of all-masked tokens ${\bm{y}}_{T}:=({\bm{M}},\ldots,{\bm{M}})$, where $L:=|{\bm{y}}_{T}|$ is a pre-specified maximum answer length. Then, in each sampling step $t\in[T]$, the MDM outputs token distributions $p_{t}^{k}:=p_{\theta}(y_{0}^{k}=\cdot\mid{\bm{x}},{\bm{y}}_{t})$ for all token positions $k\in[L]$. Then, a sampling strategy decides which subset ${\mathcal{U}}_{t}\subseteq{\mathcal{M}}_{t}$ of the still-masked positions ${\mathcal{M}}_{t}:=\{k\in[L]\mid y_{t}^{k}={\bm{M}}\}$ to unmask. A new, partially unmasked/denoised answer ${\bm{y}}_{t-1}$ is obtained via
@@
The stochasticity of sampling $y\sim p_{t}^{k}$ depends on the dLLM temperature $\tau$, with higher temperatures leading to more diverse (but often lower quality) generations. When evaluating dLLMs, it is common to set $\tau=0$ (llada2025; fastdllm2025) to maximize performance, resembling greedy decoding in AR LLMs. However, unlike for autoregressive LLMs, setting the temperature $\tau$ to 0 is not sufficient to completely determine the generation process. The choice of sampling strategy [^1] We use the terms sampling and unmasking interchangeably throughout this paper. is also important, since different choices will produce different unmasking sets ${\mathcal{U}}_{t}$. This clearly affects the sampling efficiency, since generating multiple tokens in parallel will yield a complete answer faster than if they were generated one at a time. However, it may also affect the quality of the samples, since different unmasking sets will yield different conditions ${\bm{x}}_{t}$, thus affecting the predicted token distributions $p_{t}^{k}$. The need to carefully balance efficiency and quality has thus drawn considerable attention to the development of optimal sampling techniques for dLLMs (see Section˜5).
@@ 2.2 Heuristic Samplers
Figure 1: LLaDA-8B-Instruct (llada2025) on GSM8k, with semi-AR generation (8 blocks of $BL=32$; ) and without (one block of $BL=256$; ). More datasets and models in Appendix A.1. Generation speed is measured in network function evaluations (NFEs), which corresponds to the number of sampling steps.
One popular approach to decide which positions ${\mathcal{U}}_{t}$ to unmask at each timestep is to construct a heuristic that leverages uncertainty measures derived from the predicted token distributions $p_{t}^{k}$ (fastdllm2025; ben2025accelerated; wei2025accelerating; hong2025wide; li2025beyond; huang2025pc; kim2025train; inter alia). In particular, recent work has shown substantial efficiency gains by with heuristic strategies that employ the confidence $c_{t}^{k}:=\max_{v\in{\mathcal{V}}}p_{t}^{k}(v)$ of the underlying dLLM in order to make their decisions. Two representative methods within this space are high-confidence unmasking (chang2022maskgit), in which each forward pass unmasks a fixed number of tokens ($K$) with the highest confidences,
@@
as well as the confidence-thresholding strategy of Fast-dLLM (fastdllm2025), which allows for a variable number of tokens to be unmasked at each step by comparing the confidences to a fixed threshold $\lambda$: [^2] Note that it is possible for no token confidence to exceed the pre-determined threshold $\lambda$. In this case, fastdllm2025 propose to unmask the position with the highest confidence among the remaining masked tokens, to prevent the sampling from getting stuck.
@@
Notably, fastdllm2025 showed that when combining confidence-thresholding sampling with additional optimizations such as KV-caching, dLLMs can live up to their promise and surpass AR models in token throughput, while maintaining comparable generation quality. Moreover, they provide theoretical justification for their sampling design by showing that sampling from high-confidence marginals closely approximates sampling from the joint distribution over token positions.
Despite the undeniable recent success of confidence-based samplers, relying on handcrafted solutions comes with certain drawbacks. Beyond the obvious need for careful design (e.g., selecting an appropriate confidence measure or threshold value $\lambda$), we also found that these methods often exhibit high sensitivity to the exact sampling configuration. For instance, many heuristics rely on semi-autoregressive (semi-AR) generation (arriola2025block; llada2025), where decoding proceeds one “block” of tokens at a time, with block length $BL$ serving as yet another hyperparameter. As shown in Figure 1, outside this semi-AR regime, confidence-based heuristics can degrade to below-random performance and may no longer benefit from increased compute budgets (i.e., higher numbers of function evaluations, NFEs). These limitations raise a natural question whether effective sampling methods can be learned rather than manually designed, which we explore next.
@@ 3 Learning Unmasking Policies
To facilitate the development of RL-based samplers, we start by describing the Markov decision process (MDP) (sutton1998reinforcement) for sampling in dLLMs in its most general case:  [^3] To stay aligned with the MDM notation in Section 2.1, we reverse the time in our MDP formulation: $t=T,T-1,\ldots,0$.
  • The state $({\bm{x}},{\bm{y}}_{t})$ consists of a prompt ${\bm{x}}\in{\mathcal{V}}^{d}$ and the current (partially masked) generation ${\bm{y}}_{t}\in{\mathcal{V}}^{L}$. For brevity, we omit $\bm{x}$} from the state notation unless explicitly needed. The initial state contains a fully masked generation: ${\bm{y}}_{T}=({\bm{M}},\ldots,{\bm{M}})$.
@@
  • The reward $R({\bm{y}},{\bm{y}}_{t})$ is provided only at the final generation step, which corresponds to the first timestep where all tokens have been unmasked $\hat{T}:=\max\{t\in[T]\mid y_{t}^{k}\neq M,\forall k\in[L]\}$. To learn useful samplers, the reward should promote both correctness (i.e., the generated answer $^{\bm{y}}_{\hat{T}}$ being ‘close’ to the reference answer $\bm{y}$}) and efficiency (i.e., minimizing the number of steps $T-\hat{T}$). Note that the reward depends on action ${\bm{u}}_{t}$, and therefore on the policy $\pi_{\phi}$, implicitly through its influence on the generations ${\bm{y}}_{t}$ and the number of steps; we omit it from the input arguments to simplify notation. We defer our concrete choice of the reward function to Section˜3.3.
With the MDP defined above, finding the optimal sampling becomes a standard reinforcement learning problem of finding a policy $\pi_{\phi}$ parametrized by $\phi$ that maximizes the expected reward:
$\max_{\phi}\mathbb{E}_{({\bm{x}},{\bm{y}})\sim p_{\text{data}},\>{\bm{u}}_{t}\sim\pi_{\phi},\>{\bm{y}}_{t}\sim P}\bigg[\sum_{t=0}^{T}R({\bm{y}},{\bm{y}}_{t})\bigg]\>.$

### v3

@@ Abstract
Diffusion (Large) Language Models (dLLMs) now match the downstream performance of their autoregressive counterparts on many tasks, while holding the promise of being more efficient during inference. One critical design aspect of dLLMs is the sampling procedure that selects which tokens to unmask at each diffusion step. Indeed, recent work has found that heuristic strategies such as confidence thresholding improve both sample quality and token throughput compared to random unmasking. However, such heuristics have downsides: they require manual tuning, and we observe that their performance degrades with larger block sizes. In this work, we instead propose to train sampling procedures using reinforcement learning. Specifically, we formalize masked diffusion sampling as a Markov decision process in which the dLLM serves as the environment, and propose a lightweight policy based on a single-layer transformer that maps dLLM token confidences to unmasking decisions. Our experiments show that these trained policies match the performance of state-of-the-art heuristics when combined with semi-autoregressive (block) generation, while outperforming them in the full-diffusion setting.
@@ 1 Introduction
Discrete diffusion (austin2021structured; hoogeboom2021argmax; dieleman2022continuous; DBLP:conf/icml/LouME24; DBLP:conf/nips/ShiHWDT24) has recently emerged as a compelling alternative to the predominant autoregressive (AR) modeling paradigm in large language models (LLMs). Unlike AR models, which generate the next token in a left-to-right fashion (radford2018improving), diffusion LLMs (dLLMs) generate text by learning to reverse a noising process that progressively corrupts token sequences. In particular, masked diffusion models (MDMs; sahoo2024simple; ou2024your), a subclass of discrete diffusion models that use time-dependent BERT-style (devlin2019bert) masking as the forward noising process, have recently demonstrated impressive performance, with models like LLaDA (llada2025) and Dream (ye2025dream) matching the performance of similarly sized autoregressive LLMs.
@@
At generation time, MDMs begin with a fully masked sequence and iteratively unmask a fixed number of tokens at randomly sampled positions in each sampling step. As a result, they offer the potential for faster inference, as they can, in principle, generate multiple tokens in parallel using a single model call. Despite this promise, open-source dLLMs have, until recently, lagged behind their AR counterparts in terms of inference efficiency. This changed with Fast-dLLM (fastdllm2025), which demonstrated that a dLLM like LLaDA (llada2025) can achieve higher token throughput than similarly sized LLaMA models (dubey2024llama) while maintaining competitive performance. A key component of Fast-dLLM lies in its sampling heuristic: instead of unmasking a fixed number of randomly sampled token positions, it proposes to unmask all tokens whose confidences exceed a pre-specified threshold. This success has since inspired the development of increasingly sophisticated sampling heuristics (see Appendix 5 for a comprehensive overview), further advancing the state of the art. These heuristics can, however, be difficult to tune, and seem to work best when imposing semi-autoregressive (semi-AR) generation, in which small blocks of tokens are unmasked sequentially (arriola2025block; llada2025). We posit that such limitations are difficult to avoid with heuristics alone, since they are essentially handcrafted solutions to a sequential decision making problem: What is the optimal order in which to unmask the sequence of tokens, in order to strike a good balance between correctnes and efficiency?
@@
We address this question by moving beyond heuristics, instead proposing a learning-based approach that leverages a transformer-based unmasking policy trained via reinforcement learning (RL). This approach is motivated by the above observation that unmasking in dLLMs can be viewed as a (Markovian) sequential decision making problem, and follows recent successes in RL for post-training of language models. However, in contrast to recent works that use RL to improve the reasoning abilities of dLLMs (e.g., zhao2025d1), our goal is not to improve the capacity of the underlying model, but rather to use RL to automate the discovery of adaptive sampling strategies. We therefore treat the underlying dLLM as the environment in which to act, not the policy, leaving it unchanged and instead parameterizing a policy as a lightweight stand-alone network (Figure˜2). Empirically, we demonstrate that our learned sampling policies match the performance of state-of-the-art heuristic samplers (fastdllm2025) in standard semi-AR generation, while also surpassing heuristics in full-diffusion setting. In summary, we make the following contributions:
@@
  • Our experiments show that learned policies with our RL framework match the performance of heuristic samplers such as Fast-dLLM (Section˜4.1), while also addressing some of the challenges heuristic samplers face outside the semi-AR regime (Section˜4.2).
  • We study the transferability of our learned samplers across different models and data domains (Section˜4.3), and provide detailed insights into stabilizing RL training and policy design (Section˜4.4). We also demonstrate that policies exhibit qualitatively different unmasking behavior compared to confidence heuristics (Appendix C).
@@ 2 Background
Throughout the paper, we use the notation $[L]$ to represent the set $\{1,\ldots,L\}$. We denote a sequence of tokens as ${\bm{x}}=(x^{1},\ldots,x^{d})\in{\mathcal{V}}^{d}$, where $d$ is the sequence length and ${\mathcal{V}}:=[V]$ is the vocabulary. Given our focus on MDMs, we assume the vocabulary includes a special mask token $\bm{M}$}.
@@ 2.1 Masked Diffusion Models
We focus on masked diffusion models (MDMs), deferring a more general introduction to discrete diffusion to Appendix G.
Training. MDMs learn to generate data by training a BERT-style (devlin2019bert) masked predictor to reverse the forward noising process. Concretely, given a training sample ${\bm{x}}_{0}\sim p_{\text{data}}$, the forward process corrupts each token independently by setting it to the mask token with probability proportional to the diffusion timestep $t\in[0,1]$:
$p_{t}({\bm{x}}_{t}\mid{\bm{x}}_{0}):=\prod_{k=1}^{d}p_{t}(x_{t}^{k}\mid{\bm{x}}_{0})\,,\quad\text{where }p_{t}(x_{t}^{k}\mid{\bm{x}}_{0}):=\begin{cases}1-t,&\text{if }x_{t}^{k}=x_{0}^{k},\\ t,&\text{if }x_{t}^{k}={\bm{M}},\\ 0,&\text{otherwise.}\end{cases}$
Recent work has shown that an MDM parametrized by $\theta$ can learn the reverse process by maximizing the following evidence lower bound (ELBO) (ou2024your), ${\mathcal{L}}(\theta)\leq\mathbb{E}_{{\bm{x}}_{0}\sim p_{\text{data}}}[\log p_{\theta}({\bm{x}}_{0})]$, denoting $\mathbf{1}[\cdot]$ the indicator function:
${\mathcal{L}}(\theta)$ $:=\mathbb{E}_{t\sim U[0,1],\,{\bm{x}}_{0}\sim p_{\text{data}},\,{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t}\mid{\bm{x}}_{0})}\left[\frac{1}{t}\sum_{k=1}^{d}\mathbf{1}[x_{t}^{k}={\bm{M}}]\,\log p_{\theta}(x_{0}^{k}\mid{\bm{x}}_{t})\right]\>.$ (2.1)
Generation. When generating an answer for a given prompt $\bm{x}$}, MDMs start with a sequence of all-masked tokens ${\bm{y}}_{T}:=({\bm{M}},\ldots,{\bm{M}})$, where $L:=|{\bm{y}}_{T}|$ is a pre-specified maximum answer length. Then, in each sampling step $t\in[T]$, the MDM outputs token distributions $p_{t}^{k}:=p_{\theta}(y_{0}^{k}=\cdot\mid{\bm{x}},{\bm{y}}_{t})$ for all token positions $k\in[L]$ and a sampling strategy decides which subset ${\mathcal{U}}_{t}\subseteq{\mathcal{M}}_{t}$ of the still-masked positions ${\mathcal{M}}_{t}:=\{k\in[L]\mid y_{t}^{k}={\bm{M}}\}$ to unmask. A new, partially unmasked/denoised answer ${\bm{y}}_{t-1}$ is obtained via
@@
The stochasticity of sampling $y\sim p_{t}^{k}$ depends on the dLLM temperature $\tau$, with higher temperatures leading to more diverse (but often lower quality) generations. When evaluating dLLMs, it is common to set $\tau=0$ (llada2025; fastdllm2025), resembling greedy decoding. The choice of sampling/unmasking strategy is also important, since different choices will produce different unmasking sets ${\mathcal{U}}_{t}$, thus affecting both sampling quality and efficiency. This has drawn considerable attention to the development of optimal sampling techniques for dLLMs (see Appendix 5).
@@ 2.2 Heuristic Samplers
Figure 1: LLaDA-8B-Instruct (llada2025) on GSM8k, with semi-AR generation ($BL=32$; ) and without (full-diffusion regime, $BL=L=256$; ). More datasets and models in Appendix B.1. Generation speed is measured in network function evaluations (NFEs), which corresponds to the number of sampling steps.
One popular approach to decide which positions ${\mathcal{U}}_{t}$ to unmask at each timestep is to construct a heuristic that leverages uncertainty measures derived from the predicted token distributions $p_{t}^{k}$ (fastdllm2025; ben2025accelerated; inter alia). In particular, recent work has shown substantial efficiency gains with heuristic strategies that employ the confidence $c_{t}^{k}:=\max_{v\in{\mathcal{V}}}p_{t}^{k}(v)$ of the underlying dLLM in order to make their decisions. Two representative methods within this space are high-confidence unmasking (chang2022maskgit), in which each forward pass unmasks a fixed number of tokens ($K$) with the highest confidences,
@@
as well as the confidence-thresholding strategy of Fast-dLLM (fastdllm2025), which allows for a variable number of tokens to be unmasked at each step by comparing the confidences to a fixed threshold $\lambda$:
@@
Despite the undeniable success of confidence-based samplers, relying on handcrafted solutions comes with drawbacks. Beyond the obvious need for careful design (e.g., selecting an appropriate confidence measure or threshold $\lambda$), we also found that these methods often exhibit high sensitivity to the exact sampling configuration. For instance, many heuristics rely on semi-autoregressive (semi-AR) generation (arriola2025block; llada2025), where diffusion decoding proceeds one “block” of contiguous tokens at a time, with block length $BL$ serving as yet another hyperparameter. As shown in Figure 1, outside this semi-AR regime, the performance of confidence-based heuristics can degrade below that of a random unmasking, and no longer benefit from increased compute (i.e., higher numbers of function evaluations, NFEs). These limitations raise the question of whether effective sampling methods can be learned rather than manually tuned.
@@ 3 Learning Unmasking Policies
To facilitate the development of RL-based samplers, we describe first the Markov decision process (MDP) (sutton1998reinforcement) for sampling in dLLMs. To stay aligned with the MDM notation in Section˜2.1, we reverse the time in our MDP formulation: $t=T,T-1,\ldots,1$, with $T$ denoting the maximum horizon (number of sampling steps).
  • The state $({\bm{x}},{\bm{y}}_{t})$ holds a prompt ${\bm{x}}\in{\mathcal{V}}^{d}$ and the current (partially masked) generation ${\bm{y}}_{t}\in{\mathcal{V}}^{L}$. For brevity, we omit $\bm{x}$} from the state notation unless needed. The initial state has a fully masked generation: ${\bm{y}}_{T}=({\bm{M}},\ldots,{\bm{M}})$.
@@
  • The reward $R({\bm{y}},{\bm{y}}_{t})$ is provided only at the final generation step, which corresponds to the first timestep where all tokens have been unmasked $\hat{T}:=\max\{t\in[T]\mid y_{t}^{k}\neq{\bm{M}},\forall k\in[L]\}$. To learn useful samplers, the reward should promote both correctness (i.e., the generated answer $^{\bm{y}}_{\hat{T}}$ being ‘close’ to the reference answer $\bm{y}$}) and efficiency (i.e., minimizing the number of steps $T-\hat{T}$). Note that the reward depends on action ${\bm{u}}_{t}$, and therefore on the policy $\pi_{\phi}$, implicitly through its influence on the generations ${\bm{y}}_{t}$ and the number of steps. We defer our concrete choice of the reward function to Section˜3.3.
With the MDP above, finding the optimal sampling becomes a standard RL problem of finding a policy $\pi_{\phi}$ parametrized by $\phi$ that maximizes the expected reward:
$\max_{\phi}\mathbb{E}_{({\bm{x}},{\bm{y}})\sim p_{\text{data}},\>{\bm{u}}_{t}\sim\pi_{\phi},\>{\bm{y}}_{t}\sim P}\bigg[\sum_{t=1}^{T}R({\bm{y}},{\bm{y}}_{t})\bigg]\>.$